# Rolling in the Shadows: Analyzing the Extraction of MEV Across Layer-2 Rollups

### v1

@@ 6.2. Countermeasures
Private pools [27], including encrypted mempools [34], provide an effective countermeasure against strategies $S_{1}$ and $S_{2}$ since attackers are not able to observe victim transactions via the mempool anymore and thus cannot frontrun them. However, strategy $S_{3}$ still remains feasible as it does not rely on victim transactions being visible before a block is finalized on L1. Strategy $S_{3}$ exploits the fact that there is a delay between L1 and L2 transactions. Rollups could try to reduce the delay almost to zero, such that attackers cannot react fast enough. However, this would make rollups vulnerable to time bandit attacks or uncle bandit attacks, which are already viable issues on Ethereum [31]. Another solution could be randomized ordering of transactions. However, this would break the first-come, first-served ordering policy of many rollups and would not entirely solve the issue of sandwich attacks as shown by other works [57].
@@ 7. Related Work
The first work to study MEV on rollups was by Ha et al. [30], which measured MEV on Optimism and Polygon. These measurements, however, persisted for only a couple months and did not consider Arbitrum or any ZK rollups which had not been released yet. Also at the intersection of MEV and rollups, Bagourd and Francois [9] scanned Polygon, Optimism, and Arbitrum for MEV. However, they did not scan for sandwiches, or analyze the negative aspects of MEV, nor did they compare their results to Ethereum. Besides providing results for a longer measurement period, our methodology has found more instances of MEV for the same period time than Bagourd and Francois [9] reported. Recently, Öz et al. [45] have studied MEV extraction on Algorand, which follows a first-come, first-served policy similar to rollups.
@@ Appendix A List of Used Events
Appendix B List of Identified Potential Victims
@@
Appendix C Detailed MEV Extraction Overview
Table 9 provides the total number of arbitrages, liquidations, and sandwiches extracted across Ethereum, Arbitrum, Optimism, an zkSync for our measurement period of nearly 3 years. I also provides additional granularity on the distribution of different MEV types across different DeFi protocols.
Arbitrage Liquidation Sandwiching DeFi Protocol Ethereum Arbitrum Optimism zkSync Ethereum Arbitrum Optimism zkSync Ethereum Arbitrum Optimism zkSync Aave - - - - $30734$ $2829$ $3358$ Balancer V1 $349464$ Balancer V2 $127866$ $167294$ $83044$ Compound - - - - $24038$ $2448$ $1989$ $325$ Curve $45315$ $27238$ $30186$ Uniswap V2 $2572061$ $1074085$ $990746$ $108064$ $2089610$ Uniswap V3 $1450698$ $1534733$ $830359$ $2890$ $265843$ Total Unique $2901740$ $1746083$ $1153362$ $108070$ $54772$ $5277$ $5347$ $325$ $2334566$
Table 9. Total number of arbitrages, liquidations, and sandwiches detected across different DeFi protocols and chains.
Appendix D Cross-Layer Sandwich Profits
Arbitrum Optimism Attack Strategy 1K-Cap. 10K-Cap. 100K-Cap. 1M-Cap. $\infty$-Cap. 1K-Cap. 10K-Cap. 100K-Cap. 1M-Cap. $\infty$-Cap. $S_{1}$ $P_{\text{Total}}$ 38.93 $1208.59$ $12976.01$ $122775.31$ $1215994.05$ 226.35 $4966.80$ $68991.44$ $264374.17$ $739695.63$ $P_{\text{Max}}$ 24.39 175.71 $1310.81$ $35570.84$ $785862.73$ 108.03 $1533.57$ $7481.66$ $7481.66$ $138184.20$ $P_{\text{Mean}}$ 3.89 14.92 66.89 483.37 $4374.08$ 32.34 25.34 129.20 342.00 896.60 $P_{\text{Median}}$ 0.71 7.72 29.82 82.73 105.08 6.33 8.30 50.30 165.32 187.80 $P_{\text{Min}}$ 0.00 0.00 0.00 0.00 0.00 1.31 0.07 0.41 0.41 0.41 $S_{2}$ $P_{\text{Total}}$ 59.93 $1606.38$ $15775.96$ $131257.43$ $1225968.21$ 295.37 $6688.07$ $75679.86$ $276412.81$ $752736.05$ $P_{\text{Max}}$ 27.20 178.52 $1312.94$ $35576.04$ $785866.70$ 134.46 $1560.01$ $7508.09$ $7508.09$ $138205.96$ $P_{\text{Mean}}$ 3.53 12.75 66.57 444.94 $3807.35$ 10.19 22.15 117.70 327.12 838.24 $P_{\text{Median}}$ 1.38 6.59 34.21 93.08 108.08 0.89 9.48 43.83 148.25 168.42 $P_{\text{Min}}$ 0.04 0.01 0.21 0.45 0.45 0.02 0.01 0.04 0.04 0.04 $S_{3}$ $P_{\text{Total}}$ 237.47 $2410.52$ $17108.52$ $132926.89$ $1227808.79$ $1213.84$ $11369.12$ $86369.52$ $290464.40$ $767972.37$ $P_{\text{Max}}$ 30.00 181.33 $1315.07$ $35581.23$ $785870.67$ 160.89 $1586.44$ $7534.52$ $7534.52$ $138227.72$ $P_{\text{Mean}}$ 1.10 9.49 61.76 415.40 $3558.87$ 1.71 13.95 101.97 301.00 756.62 $P_{\text{Median}}$ 0.53 4.82 27.59 84.10 101.04 0.57 4.73 37.46 119.93 134.35 $P_{\text{Min}}$ 0.00 0.04 0.04 0.04 0.04 0.00 0.00 0.04 0.04 0.04
@@ Appendix E Testnet Transactions
Table 11 provides an overview of successfully deployed transactions on the Sepolia testnet for the three attack strategies across Arbitrum, Optimism, and zkSync. Attacker transactions are highlighted in red, whereas victim transactions are highlighted in blue. These proofs-of-concept show that attackers could write scripts to automatically perform cross-layer sandwich attacks in real-time.

### v2

@@ 6.2. Countermeasures
Private pools [29], including encrypted mempools [37], provide an effective countermeasure against strategies $S_{1}$ and $S_{2}$ since attackers are not able to observe victim transactions via the mempool anymore and thus cannot frontrun them. However, strategy $S_{3}$ still remains feasible as it does not rely on victim transactions being visible before a block is finalized on L1. Strategy $S_{3}$ exploits the fact that there is a delay between L1 and L2 transactions. Rollups could try to reduce the delay almost to zero, such that attackers cannot react fast enough. However, this would make rollups vulnerable to time bandit attacks or uncle bandit attacks, which are already viable issues on Ethereum [33]. Another solution could be randomized ordering of transactions. However, this would break the first-come, first-served ordering policy of many rollups and would not entirely solve the issue of sandwich attacks as shown by previous works [61]. Finally, a number of works aim to counter sandwiching on AMMs by proposing new frontrunning resistant AMM protocols [72, 36, 63]. However, these protocols were exclusively designed to protect AMMs against frontrunning and not any other protocols. Moreover, it is not clear whether these protocols are still applicable in the context of cross-layer sandwiching.
@@ 7. Related Work
The first work to study MEV on rollups was by Ha et al. [32], which measured MEV on Optimism and Polygon. These measurements, however, persisted for only a couple months and did not consider Arbitrum or any ZK rollups which had not been released yet. Also at the intersection of MEV and rollups, Bagourd and Francois [9] scanned Polygon, Optimism, and Arbitrum for MEV. However, they did not scan for sandwiches, or analyze the negative aspects of MEV, nor did they compare their results to Ethereum. Besides providing results for a longer measurement period, our methodology has found more instances of MEV for the same time period than Bagourd and Francois [9] reported. Recently, Öz et al. [49] have studied MEV extraction on Algorand, which follows a first-come, first-served policy similar to rollups.
@@ Appendix A List of Used Events
Appendix B Detailed MEV Extraction Overview
Table 8 provides the total number of arbitrages, liquidations, and sandwiches extracted across Ethereum, Arbitrum, Optimism, an zkSync for our measurement period of nearly 3 years. It also provides additional granularity on the distribution of different MEV types across different DeFi protocols.
Appendix C List of Identified Potential Victims
@@
Appendix D Cross-Layer Sandwich Profits
Appendix E Testnet Transactions
Table 11 provides an overview of successfully deployed transactions on the Sepolia testnet for the three attack strategies across Arbitrum, Optimism, and zkSync. Attacker transactions are highlighted in red, whereas victim transactions are highlighted in blue. These transactions serve a proof-of-concept to show that attackers could easily write scripts to automatically perform cross-layer sandwich attacks in real-time.
@@ Appendix E Testnet Transactions
Arbitrum Optimism Attack Strategy 1K-Cap. 10K-Cap. 100K-Cap. 1M-Cap. $\infty$-Cap. 1K-Cap. 10K-Cap. 100K-Cap. 1M-Cap. $\infty$-Cap. $S_{1}$ $P_{\text{Total}}$ $38.93$ $1235.10$ $14631.47$ $129656.26$ $1224200.05$ $226.35$ $4966.80$ $68991.44$ $264374.17$ $739695.63$ $P_{\text{Max}}$ 24.39 175.71 $1310.81$ $35570.85$ $785862.73$ 108.03 $1533.57$ $7481.66$ $7481.66$ $138184.20$ $P_{\text{Mean}}$ 3.89 14.36 66.51 448.64 $3898.73$ 32.34 25.34 129.20 342.00 896.60 $P_{\text{Median}}$ 0.71 7.07 33.27 88.44 106.07 6.33 8.30 50.30 165.32 187.80 $P_{\text{Min}}$ 0.00 0.00 0.00 0.00 0.00 1.31 0.07 0.41 0.41 0.41 $S_{2}$ $P_{\text{Total}}$ 59.93 $1606.38$ $15775.96$ $131257.43$ $1225968.21$ 295.37 $6688.07$ $75679.86$ $276412.81$ $752736.05$ $P_{\text{Max}}$ 27.20 178.52 $1312.94$ $35576.04$ $785866.70$ 134.46 $1560.01$ $7508.09$ $7508.09$ $138205.96$ $P_{\text{Mean}}$ 3.53 12.75 66.57 444.94 $3807.35$ 10.19 22.15 117.70 327.12 838.24 $P_{\text{Median}}$ 1.38 6.59 34.21 93.08 108.08 0.89 9.48 43.83 148.25 168.42 $P_{\text{Min}}$ 0.04 0.01 0.21 0.45 0.45 0.02 0.01 0.04 0.04 0.04 $S_{3}$ $P_{\text{Total}}$ 237.47 $2410.52$ $17108.52$ $132926.89$ $1227808.79$ $1213.84$ $11369.12$ $86369.52$ $290464.40$ $767972.37$ $P_{\text{Max}}$ 30.00 181.33 $1315.07$ $35581.23$ $785870.67$ 160.89 $1586.44$ $7534.52$ $7534.52$ $138227.72$ $P_{\text{Mean}}$ 1.10 9.49 61.76 415.40 $3558.87$ 1.71 13.95 101.97 301.00 756.62 $P_{\text{Median}}$ 0.53 4.82 27.59 84.10 101.04 0.57 4.73 37.46 119.93 134.35 $P_{\text{Min}}$ 0.00 0.04 0.04 0.04 0.04 0.00 0.00 0.04 0.04 0.04